# Democratizing Signal Processing and Machine Learning: Math Learning Equity for Elementary and Middle School Students

### v1

@@ III Partial solution 1: out-of-school math programs such as CyMath and Ab7G
  • Practice in school. Our first suggestion is that elementary schools make students do enough math practice during school itself. This would be easy for teachers to incorporate if they do not have to convert every math lesson into a game. Planning for fun math games for each topic takes a lot of the teacher’s time and uses up a lot of classroom time as well. It likely takes away the class time that could otherwise be spent by the students on math practice; and by the teachers on checking the students’ work and correcting math mistakes early. In fact, reducing the time and effort needed to make all learning fun may also leave the teacher with more time for one-on-one support for students who need that.
  • Homework. Our second suggestion is that schools send reasonable amounts of math work home and require students to complete it. Individual math practice is essential for math fluency. Moreover, teachers should be allowed time to check students’ work carefully so that student learning can be corrected early.
@@
  • Awareness. Most parents want the information needed to help their child learn and do better in school and later life. However, many parents are unaware that math learning is cumulative and hence a good early foundation in math is important, or that there are two or more math tracks starting in fifth or sixth grade in many countries. They also do not know how to help their child learn math better. This knowledge should not be the privilege of STEM professional parents or their relatives. It should be provided to everyone by the schools. Moreover, it will automatically get conveyed if math is emphasised, homework is sent, and free workbooks, printed worksheets, and information about good online resources such as Khan academy, are shared with families. These cost almost nothing. Math should be the least expensive subject to improve outcomes in, since it requires no equipment or supplies. Some information is summarized here Math-info.
@@
  • Math tests awareness and feedback. This last suggestion may apply to the US only. There are few regular low-stakes tests in elementary schools, and, even when there are, very little feedback is provided to parents based on the test results. At the same time, schools do use state-wide tests, such as the Measure of Academic Performance (MAP) test, given a few times a year to evaluate students and to place them in different math tracks. However parents are not made sufficiently aware of this fact ahead of time. Students, and parents, are often not sufficiently informed about the test dates or that the tests are adaptive (and what that means). This is done in order to not put any demands on parents’ time, and follows advice provided by published literature. At the same time, if a parent checks with the teacher, they are provided all the information. Consequently, we end up disadvantaging a much larger number of children who may have otherwise come better prepared, or at least better rested, for these tests. Good test taking skills, and studying for a test, are also important for college success.
@@
  • Use of summer to make-up for lost learning; relying on STEM students and professionals. Despite all efforts, some students may require additional support to stay on track. There should also be easily available literature, and resources, for helping students catch-up with lost learning at various stages. For example, for a pre-algebra student with learning gaps, as long as they can learn at least one way of doing all the arithmetic, middle school algebra will make sense to them. The summer before middle school can be used efficiently to help many students catch up. Summer is also the time many undergraduate and graduate students and faculty are more available and schools can tap into these resources by partnering with nearly colleges or universities.

### v2

@@ IV Lessons learned and Conclusions
    Besides homework, providing enough in-school individual math practice can also help with student learning in younger grades. This would be easier for teachers to incorporate if they were not encouraged convert every math lesson into a game. Planning for fun math games takes a lot of teacher time and uses up a lot of classroom time. It likely takes away the class time that could otherwise be spent by the students on math practice; and by the teachers on checking the students’ work and correcting errors early.
@@
  • Parent Awareness. Many parents are not aware that a good early foundation in math is important, or how to help their child build such a foundation. This knowledge should not be the privilege of children or relatives of STEM professionals. It should be provided to everyone by the schools. Moreover, it will automatically get conveyed if math is emphasised, homework is sent, and if free workbooks, printed worksheets, and information about good online resources such as Khan academy or k5learning, are shared with families. Some information is summarized here Math-info.
@@
    Moreover, in many elementary schools, there are few tests, and little feedback is provided to parents based on the test results. Parents or students are often not even informed about the test dates even though the same tests’ results are used later to place students into different math tracks. However, if a parent asks, they are provided all of this information. Consequently, we end up disadvantaging a much larger number of children who may have otherwise come better prepared for these tests. Good test taking skills are also quite important for college success.
@@
  • Use of summer to make-up for lost learning. Despite all efforts, some students may require additional support to stay on track. The summer before middle school can be used efficiently to help students catch up on arithmetic and pre-algebra skills for example. Summer is also the time many undergraduate and graduate students and faculty are more available and schools can tap into these resources by partnering with nearly colleges or universities.